# Physical Intelligence

### v1

@@ V-A Pre-training and post-training
Figure 4: Overview of our dataset: The pre-training mixture consists of a subset of OXE [10] and the $\pi$ dataset. We use a subset of OXE, which we refer to as OXE Magic Soup [24]. The left figures illustrates the weight of the different datasets in the pre-training mixture. The right figure illustrates their relative sizes as measured by the number of steps.
@@ VI-A Evaluating the base model
Figure 7: Zero-shot evaluation results: We evaluate $\pi_{0}$ trained for the full 700k steps, a version trained for 160k steps that matches the number of updates for baseline models, $\pi_{0}$-small, and three baselines: OpenVLA and Octo trained on all of our data, and OpenVLA trained only on the UR5e tasks (which we found to work a bit better on these tasks). Across all tasks and all comparisons, even the “parity” version of our model outperforms all baselines, and the full version of our model achieves the best results by a large margin.
@@
Bussing hard: a harder version of the bussing task, with more objects and more challenging configurations, such as utensils intentionally placed on top of trash objects and some objects that are not in the pre-training dataset.
@@ VI-B Following language commands
Figure 9: Language evaluation. We compare “flat” versions of our policies, $\mathtt{-flat}$, which receive only the overall task command (e.g., “bag the groceries”) with a method that receives intermediate commands from a human expert, $\mathtt{-human}$, or a high-level VLM policy, $\mathtt{-HL}$. We also compare our model to a small non-VLM variant under the “expert” condition, $\pi_{0}$ and $\pi_{0}$-small, in terms of language following accuracy. The results show a significant improvement with $\pi_{0}$ from intermediate language commands provided by a human expert and to a lesser degree by an autonomous high-level policy. Crucially, due to $\pi_{0}$-small’s language following performance, it does not gain with the addition of a high-level expert.
@@
The results in Figure 9, averaging over 10 trials per task, show that the language following accuracy of $\pi_{0}$ is significantly better than that of $\pi_{0}$-small. This suggests a significant improvement from the larger pre-trained VLM initialization. This capability translates to an improvement in performance with expert human guidance ($\pi_{0}$-human) and with high-level model guidance ($\pi_{0}$-HL). The results indicate that our model’s language following ability directly translates into better autonomous performance on complex tasks with high-level guidance.
@@ VI-C Learning new dexterous tasks
Figure 11 shows the performance across all of the tasks for a variety of methods, averaging over 10 trials per task, with different amounts of fine-tuning data on each task. We include all of the baselines on the stack bowls and tupperware in microwave tasks. Since OpenVLA and Octo attain significantly worse performance, we only run these for one the dataset sizes, due to the time cost of evaluating so many models in the real world. The results show that $\pi_{0}$ generally outperforms other methods. Interestingly, the strongest prior models are the ones that are trained entirely from scratch on the target tasks, suggesting that leveraging pre-training in these domains presents a major challenge for prior approaches. While the 5-hour policy for $\pi_{0}$ on the tuppaware task performs similarly to the baselines, the 1-hour version is significantly better. As expected, pre-training leads to larger improvement for tasks that are more similar to the pre-training data, though the pre-trained model is frequently better than the non-pre-trained model, sometimes by as much as 2x.
@@ VI-D Mastering complex multi-stage tasks
Figure 13: Post-training results on complex tasks in terms of average scores over 10 trials. The full pre-trained $\pi_{0}$ model attains more than 50% of the maximum score across all of the tasks, and typically outperforms the ablations, with esepcially significant improvements on the hardest tasks.
@@
Mobile dryer: Here, the Fibocom mobile robot has to take laundry out of a dryer and place it into a hamper. This task is present in pre-training.
@@ A-A Contributions
Data: Noah Brown, Michael Equi, Chelsea Finn, Niccolo Fusai, Lachy Groom, Liyiming Ke, Suraj Nair, Lucy Shi, and Anna Walling.
Evaluation experiments: Chelsea Finn, Michael Equi, Brian Ichter, Liyiming Ke, Adrian Li-Bell, Suraj Nair, Karl Pertsch, and Lucy Shi.
@@
Post-training: Chelsea Finn, Michael Equi, Liyiming Ke, Adrian Li-Bell, Suraj Nair, and Lucy Shi.
@@ A-D Inference
Recall that our model takes an observation $\mathbf{o}_{t}=[\mathbf{I}^{1}_{t},...,\mathbf{I}^{n}_{t},\ell_{t},\mathbf{q}_{t}]$ and the noisy actions $\mathbf{A}_{t}^{\tau}$ and outputs the vector field that needs to be integrated to obtain the next flow matching step, $\mathbf{v}_{t}^{\tau}$. Each time we predict a new action chunk $\mathbf{A}_{t}$, we must encode each of the images $\mathbf{I}^{1}_{t},...,\mathbf{I}^{n}_{t}$, run a forward pass on the tokens corresponding to $\mathbf{o}_{t}$, and then run 10 steps of flow matching, where each step requires running a forward pass on the tokens corresponding to $\mathbf{A}_{t}^{\tau}$ (the keys and values corresponding to $\mathbf{o}_{t}$ are cached). Table I summarizes the computation time for this operation with 3 camera images. The operations were timed on an NVIDIA GeForce RTX 4090 consumer-grade GPU. For the mobile robot, inference was done off-board over a WiFi connection, adding a small amount of network latency. Further optimizations, quantization, and other improvements might further reduce inference times.
@@
TABLE I: Inference time of our model on a NVIDIA GeForce RTX 4090 GPU.
@@ A. Evaluating the base model
Grocery bagging: This task is scored out of 7. For each 7 grocery items, a poinr is given for putting it in the bag.
@@
Grocery bagging: The robot picks up the correct item (among bag of coffee beans, bag of barley, bag of marshmallow, cat can food, spaghetti, bag of seaweed, bag of almonds), and bags them into a paper bag. The robot receives 7 objects in total and around 14 instructions in one episode.
@@ D. Mastering complex multi-stage tasks
Laundry folding: This task is scored out of 4. Our evaluation includes five items, three shirts of size M, L, and XL and two shorts of size 28 and 36. We perform two trials for each item, and the items left to be evaluated start randomly crumpled in a laundry bin (while previously evaluted items start in a folded stack). One point is given for picking an item out of the bin and putting it on the table. Another point is given for flattening the shirt or shorts. A third point is granted for folding the shirt or shorts. A final point is given for either placing the item in the corner of the table (if it is the first item evaluated), or stacking it onto an existing stack of folded clothes. We run each eval for a maximum of 15000 steps or approximately 5 minutes.
@@
Box building: This task is scored out of 5. One point is given for successfully picking up the box to begin the task. One point is given for folding the box in half so the flaps can be closed. One point is given for closing the right flap. One point is given for closing the left flap. The final point is given for neatly centering the final product.
@@
Dryer unloading: This task involves having the robot approach a dyer with a laundry basket and unload the clothes into the basket. We score this eval out of five where one point is given for properly approaching the dryer. Another for placing the laundry basket on the stool. A third for opening the dryer. A fourth for putting all the clothes in the basket and a fifth point for closing the dryer. We eval with 3 shirts and 2 shorts that start in a random configuration inside the dryer.

### v2

@@ V-A Pre-training and post-training
Figure 4: Overview of our dataset: The pre-training mixture consists of a subset of OXE [10] and the $\pi$ dataset. We use a subset of OXE, which we refer to as OXE Magic Soup [24]. The right figure illustrates the weight of the different datasets in the pre-training mixture. The left figure illustrates their relative sizes as measured by the number of steps.
@@ VI-A Evaluating the base model
Figure 7: Zero-shot evaluation results: We evaluate $\pi_{0}$ trained for the full 700k steps, a version trained for 160k steps that matches the number of updates for baseline models, $\pi_{0}$-small, and three baselines: OpenVLA and Octo trained on all of our data, and OpenVLA trained only on the UR5e tasks (which we found to work better on UR5e tasks). Across all tasks and all comparisons, even the “parity” version of our model outperforms all baselines, and the full version of our model achieves the best results by a large margin.
@@
Bussing hard: a harder version of the bussing task, with more objects and more challenging configurations, such as utensils intentionally placed on top of trash objects, objects obstructing each other, and some objects that are not in the pre-training dataset.
@@ VI-B Following language commands
Figure 9: Language evaluation. We compare “flat” versions of our policies, $\mathtt{-flat}$, which receive only the overall task command (e.g., “bag the groceries”) with a method that receives intermediate commands from a human expert, $\mathtt{-human}$, or a high-level VLM policy, $\mathtt{-HL}$. We also compare our model to a small non-VLM variant under the “expert” condition, $\pi_{0}$ and $\pi_{0}$-small, in terms of language following accuracy. The results show a significant improvement with $\pi_{0}$ from intermediate language commands provided by a human expert and to a lesser degree by an autonomous high-level policy. Notably, due to $\pi_{0}$-small’s limited language following ability, overall it does not gain with the addition of a high-level expert.
@@
The results in Figure 9, averaging over 10 trials per task, show that the language following accuracy of $\pi_{0}$ is significantly better than that of $\pi_{0}$-small. This suggests a significant improvement from the larger pre-trained VLM initialization. This capability translates to an improvement in performance with expert human guidance ($\pi_{0}$-human) and with high-level model guidance ($\pi_{0}$-HL). The results indicate that $\pi_{0}$’s language following ability directly translates into better autonomous performance on complex tasks with high-level guidance.
@@ VI-C Learning new dexterous tasks
Figure 11 shows the performance across all of the tasks for a variety of methods, averaging over 10 trials per task, with different amounts of fine-tuning data on each task. We include all of the baselines on the stack bowls and Tupperware in microwave tasks. Since OpenVLA and Octo attain significantly worse performance, we only run these for one of the dataset sizes, due to the time cost of evaluating so many models in the real world. The results show that $\pi_{0}$ generally outperforms other methods. Interestingly, the strongest prior models are the ones that are trained entirely from scratch on the target tasks, suggesting that leveraging pre-training in these domains presents a major challenge for prior approaches. While the 5-hour policy for $\pi_{0}$ on the Tupperware task performs similarly to the baselines, the 1-hour version is significantly better. As expected, pre-training leads to larger improvement for tasks that are more similar to the pre-training data, though the pre-trained model is frequently better than the non-pre-trained model, sometimes by as much as 2x.
@@ VI-D Mastering complex multi-stage tasks
Figure 13: Post-training results on complex tasks in terms of average scores over 10 trials. The full pre-trained $\pi_{0}$ model attains more than 50% of the maximum score across all of the tasks, and typically outperforms the ablations, with especially significant improvements on the hardest tasks.
@@
Dryer unloading: Here, the Fibocom mobile robot has to take laundry out of a dryer and place it into a hamper. This task is present in pre-training.
@@ A-A Contributions
Data and operations: Noah Brown, Michael Equi, Chelsea Finn, Niccolo Fusai, Lachy Groom, Liyiming Ke, Suraj Nair, Lucy Shi, and Anna Walling.
Evaluation experiments: Kevin Black, Michael Equi, Chelsea Finn, Brian Ichter, Liyiming Ke, Adrian Li-Bell, Suraj Nair, Karl Pertsch, and Lucy Shi.
@@
Post-training: Michael Equi, Chelsea Finn, Liyiming Ke, Adrian Li-Bell, Suraj Nair, and Lucy Shi.
@@ A-D Inference
Recall that our model takes an observation $\mathbf{o}_{t}=[\mathbf{I}^{1}_{t},...,\mathbf{I}^{n}_{t},\ell_{t},\mathbf{q}_{t}]$ and the noisy actions $\mathbf{A}_{t}^{\tau}$ and outputs the vector field that needs to be integrated to obtain the next flow matching step, $\mathbf{v}_{t}^{\tau}$. Each time we predict a new action chunk $\mathbf{A}_{t}$, we must encode each of the images $\mathbf{I}^{1}_{t},...,\mathbf{I}^{n}_{t}$, run a forward pass on the tokens corresponding to $\mathbf{o}_{t}$, and then run 10 steps of flow matching, where each step requires running a forward pass on the tokens corresponding to $\mathbf{A}_{t}^{\tau}$ (the keys and values corresponding to $\mathbf{o}_{t}$ are cached). Table I summarizes the computation time for this operation with 3 camera images. The operations were timed on an NVIDIA GeForce RTX 4090 consumer-grade GPU. For the mobile robot, inference was done off-board over a Wi-Fi connection, adding a small amount of network latency. Further optimizations, quantization, and other improvements might further reduce inference times.
@@
TABLE I: Inference time of our model on an NVIDIA GeForce RTX 4090 GPU.
@@ A. Evaluating the base model
Grocery bagging: This task is scored out of 7. For each 7 grocery items, a point is given for putting it in the bag.
@@
Grocery bagging: The robot picks up the correct item (among bag of coffee beans, bag of barley, bag of marshmallow, cat food, spaghetti, bag of seaweed, bag of almonds), and bags them into a paper bag. The robot receives 7 objects in total and around 14 instructions in one episode.
@@ D. Mastering complex multi-stage tasks
Laundry folding: This task is scored out of 4. Our evaluation includes five items, three shirts of size M, L, and XL and two shorts of size 28 and 36. We perform two trials for each item, and the items left to be evaluated start randomly crumpled in a laundry bin (while previously evaluated items start in a folded stack). One point is given for picking an item out of the bin and putting it on the table. Another point is given for flattening the shirt or shorts. A third point is granted for folding the shirt or shorts. A final point is given for either placing the item in the corner of the table (if it is the first item evaluated), or stacking it onto an existing stack of folded clothes. We run each eval for a maximum of 15000 steps or approximately 5 minutes.
@@
Box building: This task is scored out of 5. One point is given for successfully picking up the box to begin the task. One point is given for folding the box in half, so the flaps can be closed. One point is given for closing the right flap. One point is given for closing the left flap. The final point is given for neatly centering the final product.
@@
Dryer unloading: This task involves having the robot approach a dyer with a laundry basket and unload the clothes into the basket. We score this eval out of five, where one point is given for properly approaching the dryer. Another for placing the laundry basket on the stool. A third for opening the dryer. A fourth for putting all the clothes in the basket and a fifth point for closing the dryer. We eval with 3 shirts and 2 shorts that start in a random configuration inside the dryer.